# Dynamic and Thermodynamic Stability of Superconducting-superfluid Stars

### v1

@@ III Lagrangian framework for diffeomorphism covariant theories
Finally, let us return to Eq. (43) in the case where $\phi$ has a time translation symmetry, i.e. $\mathcal{L}_{t}\phi=0$ for a vector field $t^{a}$ that approaches a time translation at infinity. We further assume that the equations of motion, $\bm{E}=0$, hold in a neighborhood of infinity, but we do not assume that they hold in the interior of the spacetime. We similarly assume that $\delta\phi$ satisfies the linearized constraints near infinity, but do not assume that these hold in the interior of the spacetime, nor do we make any symmetry assumptions on $\delta\phi$. Then the left side of Eq. (43) vanishes if $X^{a}=t^{a}$, and the surface integral on the right side simply yields $\delta M$. Thus we obtain
$\delta M=-\int_{\Sigma}\left(\iota_{X}\bm{E}\cdot\delta\phi+\delta\bm{C}_{X}\right).$ (53)
@@ IV.1 First law of thermodynamics
the constraint 3-form $\bm{C}_{X}$ is
$\left(\bm{C}_{X}\right)_{abc}=-X^{d}\bm{\epsilon}_{eabc}\left[2\left(E_{G}\right)_{\ d}^{e}+E_{F}^{e}A_{d}\right],$ (57)
and the Noether charge 2-form $\bm{Q}_{X}$ is
$\left(\bm{Q}_{X}\right)_{ab}=-\bm{\epsilon}_{abcd}\left(\nabla^{c}X^{d}+\frac{1}{2}F^{cd}X^{e}A_{e}\right).$ (58)
@@ IV.2 Existence of desired solutions to the linearized constraints
Let $\delta\mathcal{N}_{\text{X}}$, $\delta\mathcal{S}$, and $\delta\mathcal{J}_{\text{X}}$ be specified arbitrary as smooth, axisymmetric functions with support inside the background star, such that $\delta\mathcal{J}_{\text{X}}/\varphi^{a}\varphi_{a}$ also is smooth. Then we can choose the remaining initial data $\left(\delta h_{ij},\delta\pi^{ij},\delta A_{i},\delta\Pi^{i},\delta\mu_{\perp}^{\text{X}i}\right)$ so as to solve the linearized constraints Eqs. (82), (84), (85), and (86).
@@
which has a unique axisymmetric solution that goes to zero at infinity given any prescribed perturbations $\delta\mathcal{N}_{\text{X}}$ [Shi:2021dvd]. Furthermore, Eq. (84) can be cast into
@@
Since the right side is a smooth vector field of compact support, there also exists a unique solution to Eq. (90) that goes to zero at infinity [Chrusciel:2003sr].
@@
The positive definiteness of the inertia matrix Eq. (9) implies that $\mathscr{M}$ is non-negative, and since the right side of Eq. (91) vanishes suitably rapidly at infinity, then there exists a unique solution, $\psi$, of this equation that vanishes at infinity [Chrusciel:2003sr].
@@ IV.3 Thermodynamic equilibrium
If the star is in weak thermodynamic equilibrium, according to the Lemma 1, consider an arbitrary perturbation $\delta_{1}$ such that $\delta_{1}N^{\text{X}}=\delta_{1}J^{\text{X}}=0$ and $\delta_{1}S\neq 0$, then there must be $\delta_{1}M\neq 0$. Suppose that at least one of $\tilde{\mu}_{\text{X}}$, $\tilde{T}$, and $\Omega_{\text{X}}$ were not uniform throughout the star, without loss of generality, say $\tilde{T}$ is not uniform. Let
$\overline{T}=\frac{\int_{\Gamma}\tilde{T}\nu^{a}\bm{\epsilon}_{abcd}}{\int_{\Gamma}\nu^{a}\bm{\epsilon}_{abcd}},$ (95)
where $\Gamma$ is the compact support of the star. Now choose a second perturbation $\delta_{2}$ with perturbative parameter $\varepsilon$ satisfying the linearized constraint equations given by
@@
$\delta_{2}\bm{S}|_{\Gamma}$ $=\varepsilon\left(\tilde{T}-\overline{T}\right)\nu^{a}\bm{\epsilon}_{abcd},\quad\delta_{2}\bm{S}|_{\Sigma\backslash\Gamma}=0,$ (96)
we can always do this by using Lemma 1 again. With such a perturbation, one has
@@
$\delta_{2}M$ $=\int_{\Sigma}\left(\sum_{\text{X}}\tilde{\mu}_{\text{X}}\delta\bm{N}^{\text{X}}+\tilde{T}\delta\bm{S}+\sum_{\text{X}}\Omega_{\text{X}}\delta\bm{J}^{\text{X}}\right)$
$=\varepsilon\int_{\Gamma}\left(\tilde{T}-\overline{T}\right)^{2}\nu^{a}\bm{\epsilon}_{abcd}.$ (98)
@@
$\delta_{2}\bm{J}^{\text{n}}|_{\Gamma}$ $=-\delta_{2}\bm{J}^{\text{p}}|_{\Gamma}=\varepsilon\left(\Omega_{\text{n}}-\Omega_{\text{p}}\right)\nu^{a}\bm{\epsilon}_{abcd},$
$\delta_{2}\bm{J}^{\text{n}}|_{\Sigma\backslash\Gamma}$ $=\delta_{2}\bm{J}^{\text{p}}|_{\Sigma\backslash\Gamma}=0,$ (99)
@@
$\delta N=\delta S=\delta J=0,$ (100)
@@
$\delta_{2}M$ $=\int_{\Sigma}\left(\sum_{\text{X}}\tilde{\mu}_{\text{X}}\delta\bm{N}^{\text{X}}+\tilde{T}\delta\bm{S}+\sum_{\text{X}}\Omega_{\text{X}}\delta\bm{J}^{\text{X}}\right)$
$=\varepsilon\int_{\Gamma}\left(\Omega_{\text{n}}-\Omega_{\text{p}}\right)^{2}\nu^{a}\bm{\epsilon}_{abcd}.$ (101)
@@
And the equation of motion is given by [Carter:1987qr]
@@
Now, in the transfusive case, without the conserved quantities given in Eq. (74), our definition of weak thermodynamic equilibrium obviously can not work, but the definition of strong thermodynamic equilibrium still works due to the conservation of the total particle number $N$, the total entropy $S$, and the total angular momentum $J$. As shown in [Langlois:1997bz], if there is a differential rotation, the different constituents will drag each other until they have same angular velocity. This statement actually coincides our condition that there will be no differential rotation as stated in Theorem 3. On the other hand, although the three fluids with differential rotation will finally be locked together, this progress likely takes many hundreds of dynamical timescales, so when the timescale under consideration is not too long, the transfusive effect is not important, and the non-transfusive model will be a good approximation to describe the neutron star.
@@
We want to construct the phase space of our Einstein-superconducting-superfluid system. To achieve this, we will give the pre-symplectic form $W$ by using Wald formalism introduced in Sec. III, then we seek the degeneracy of $W$ and construct the phase space by factoring these degeneracy. Since in the next section, it will be important to determine the symplectic complement corresponding to field variations of the trivial perturbation within the subspace consisting of weak solutions to the linearized constraints, so we will talk about the trivial perturbations generated by the trivial displacements in the last subsection.
@@ V.1 Lagrangian and symplectic form
And the constraint 3-form $\bm{C}_{X}$ is still
$\left(\bm{C}_{X}\right)_{abc}=-X^{d}\bm{\epsilon}_{eabc}\left[2\left(E_{G}\right)_{\ d}^{e}+E_{F}^{e}A_{d}\right].$ (115)
@@
Since when the Lie derivative acting on the forms, one has $\left[\mathscr{L}_{X},\mathscr{L}_{Y}\right]=\mathscr{L}_{\left[X,Y\right]}$, so that
@@
where we have used the well known expression [Burnett:1990tww] for the gravitational and electromagnetic parts of symplectic current, and
@@
V.2 Phase space
@@
$W\left(\delta_{1}\phi,\delta_{2}\phi\right)=$ $W\left(\delta_{1}\phi,\delta_{2}^{\prime}\phi\right)+W\left(\delta_{1}\phi,\delta_{2}^{\text{M}}\phi\right)$
$=$ $\int_{\Sigma}\left(\delta_{1}\bm{\pi}^{ij}\delta_{2}h_{ij}-\delta_{2}\bm{\pi}^{ij}\delta_{1}h_{ij}\right)+\int_{\Sigma}\left(\delta_{1}\bm{\Pi}^{i}\delta_{2}A_{i}-\delta_{2}\bm{\Pi}^{i}\delta_{1}A_{i}\right)-\int_{\Sigma}\sum_{\text{X}}\xi_{\text{X}}^{a}\delta_{2}^{\prime}\bm{P}_{a}^{\text{X}}$
@@
Note that the first line only occur the quantities $\delta_{2}h_{ij}$, $\delta_{2}\bm{\pi}^{ij}$, $\delta_{2}\alpha$, $\delta_{2}\beta_{a}$, $\delta_{2}A_{a}$, $\delta_{2}\bm{\Pi}^{i}$, while the second line only occurs $\zeta_{\text{X}}^{a}$ and its normal derivative. According to
@@
where with $q_{\text{XY}}^{ab}=g^{ab}+\frac{1}{x_{\text{XY}}^{2}}n_{\text{X}}^{(a}n_{\text{Y}}^{b)}$,
@@
into Eq. (246). And also
@@
i.e., $\xi_{\text{X}}^{a}$ is the linear combination of $\nu^{a}$ and $\varphi^{a}$. Thus Eq. (147) amounts to say that the only choice is $\xi_{\text{X}}^{a}\propto u_{\text{X}}^{a}$, and substituting return to Eq. (139), one can easily see that $\xi_{\text{X}}^{a}\propto u_{\text{X}}^{a}$ is actually the solution. It immediately that Eqs. (132) and (133) are satisfied, in addition, Eqs. (134) and (137) yield $\delta_{1}\bm{\pi}^{ij}=0$ and $\delta_{1}\bm{\Pi}^{i}=0$ respectively.
@@
where we use $\cdot$” denotes the contraction of all tensor indices after using the background metric $h_{ab}$ on $\Sigma$ to raise and lower indices. Thus the elements of $\mathcal{H}$ are the square integrable tensor fields $\left(q^{\alpha},p_{\alpha}\right)$ on $\Sigma$. Note that the perturbations for which $\delta M\neq 0$ fall off too slowly to be square integrable, but $\mathcal{H}$ contains all perturbations of interest for which $M=0$.
@@
Now let $\phi$ satisfy the equations of motion and let $X^{a}$ be smooth and of compact support. By the fundamental identity Eq. (41), we have for all $\delta\phi\in\mathcal{H}$,
$\braket{\delta\phi,\hat{W}\mathscr{L}_{X}\phi}=\int_{\Sigma}X^{a}\delta\bm{C}_{a}.$ (166)
By definition, the right side vanishes if and only if $\delta\phi$ is a weak solution of the constraint equations, $\delta\bm{C}_{a}=0$. Thus if we take $\mathcal{G}$ to be subspace of $\mathcal{H}$ spanned by perturbations of the form $\mathscr{L}_{X}\phi$, we see that $\mathcal{G}^{\perp_{s}}$ is precisely the subspace, $\mathcal{C}$, of weak solutions to the linearized constraints. Furthermore, by the general argument of the previous paragraph, we have $\mathcal{C}^{\perp_{s}}=\overline{\mathcal{G}}$. Another way of saying this is that if we restrict the action of the original quadratic form to $\mathcal{C}\times\mathcal{C}$, it becomes degenerate precisely on the gauge transformations $\mathscr{L}_{X}\phi$.
V.3 Trivial displacements

### v2

@@ III Lagrangian framework for diffeomorphism covariant theories
Finally, let us return to Eq. (44) in the case where $\phi$ has a time translation symmetry, i.e. $\mathcal{L}_{t}\phi=0$ for a vector field $t^{a}$ that approaches a time translation at infinity. We further assume that the equations of motion, $\bm{E}=0$, hold in a neighborhood of infinity, but we do not assume that they hold in the interior of the spacetime. We similarly assume that $\delta\phi$ satisfies the linearized constraints near infinity, but do not assume that these hold in the interior of the spacetime, nor do we make any symmetry assumptions on $\delta\phi$. Then the left side of Eq. (44) vanishes if $\mathcal{X}^{a}=t^{a}$, and the surface integral on the right side simply yields $\delta M$. Thus we obtain
$\delta M=-\int_{\Sigma}\left(\iota_{t}\bm{E}\cdot\delta\phi+\delta\bm{C}_{t}\right).$ (54)
@@ IV.1 First law of thermodynamics
the constraint 3-form $\bm{C}_{\mathcal{X}}$ is
$\left(\bm{C}_{\mathcal{X}}\right)_{abc}=-\mathcal{X}^{d}\bm{\epsilon}_{eabc}\left[2\left(E_{G}\right)_{\ d}^{e}+E_{F}^{e}A_{d}\right],$ (58)
and the Noether charge 2-form $\bm{Q}_{\mathcal{X}}$ is
$\left(\bm{Q}_{\mathcal{X}}\right)_{ab}=-\bm{\epsilon}_{abcd}\left(\nabla^{c}\mathcal{X}^{d}+\frac{1}{2}F^{cd}\mathcal{X}^{e}A_{e}\right).$ (59)
@@ IV.2 Existence of desired solutions to the linearized constraints
Let $\delta\mathcal{N}_{\text{X}}$, $\delta\mathcal{S}$, and $\delta\mathcal{J}_{\text{X}}$ be specified arbitrary smooth, axisymmetric functions with compact support, such that $\delta\mathcal{J}_{\text{X}}/\varphi^{a}\varphi_{a}$ also is smooth. Then we can choose the remaining initial data $\left(\delta h_{ij},\delta\pi^{ij},\delta A_{i},\delta\Pi^{i},\delta\mu_{\perp}^{\text{X}i}\right)$ so as to solve the linearized constraints Eqs. (83), (85), (86), and (87).
@@
which has a unique axisymmetric solution that goes to zero at infinity given any prescribed perturbations $\delta\mathcal{N}_{\text{X}}$ [33]. Furthermore, Eq. (85) can be cast into
@@
Since the right side is a smooth vector field of compact support, there also exists a unique solution to Eq. (91) that goes to zero at infinity [17].
@@
The positive definiteness of the inertia matrix Eq. (9) implies that $\mathscr{M}$ is non-negative, and since the right side of Eq. (92) vanishes suitably rapidly at infinity, then there exists a unique solution, $\psi$, of this equation that vanishes at infinity [17].
@@ IV.3 Thermodynamic equilibrium
According to Lemma 1, consider a star in weak thermodynamic equilibrium and an arbitrary perturbation $\delta_{1}$ satisfying $\delta_{1}N^{\text{X}}=\delta_{1}J^{\text{X}}=0$ and $\delta_{1}S\neq 0$, then there must be $\delta_{1}M\neq 0$. Suppose that at least one of $\tilde{\mu}_{\text{X}}$, $\tilde{T}$, and $\Omega_{\text{X}}$ were not uniform throughout the star, without loss of generality, say $\tilde{T}$ is not uniform. As our star is a compact object, let $\Gamma$ be the compact support of the star, which is contained in some open subset $\mathcal{U}$ of $\Sigma$ whose closure is also compact. According to the existence of $C_{\infty}$ partition of unity, there exists a smooth bump function $\kappa:\Sigma\rightarrow\left[0,1\right]$ such that $\kappa\equiv 1$ on $\Gamma$ and $\text{supp}\kappa\subset\mathcal{U}$. Let
$\overline{T}=\frac{\int_{\mathcal{U}}\kappa\tilde{T}\nu^{a}\bm{\epsilon}_{abcd}}{\int_{\mathcal{U}}\kappa\nu^{a}\bm{\epsilon}_{abcd}}.$ (96)
Now choose a second perturbation $\delta_{2}$ with perturbative parameter $\varepsilon$ satisfying the linearized constraint equations given by
@@
$\delta_{2}\bm{S}|_{\mathcal{U}}$ $=\varepsilon\kappa\left(\tilde{T}-\overline{T}\right)\nu^{a}\bm{\epsilon}_{abcd},\quad\delta_{2}\bm{S}|_{\Sigma\backslash\mathcal{U}}=0,$ (97)
since clearly $\delta_{2}\bm{N}^{\text{X}}$, $\delta_{2}\bm{J}^{\text{X}}$, and $\delta_{2}\bm{S}$ are smooth with compact supports, we can always do this choice by using Lemma 1 again. With such a perturbation, one has
@@
$\delta_{2}M$ $=\int_{\Sigma}\left(\sum_{\text{X}}\tilde{\mu}_{\text{X}}\delta_{2}\bm{N}^{\text{X}}+\tilde{T}\delta_{2}\bm{S}+\sum_{\text{X}}\Omega_{\text{X}}\delta_{2}\bm{J}^{\text{X}}\right)$
$=\int_{\Sigma}\left(\tilde{T}-\overline{T}\right)\delta_{2}\bm{S}+\overline{T}\delta_{2}S$ (99)
$=\varepsilon\int_{\mathcal{U}}\kappa\left(\tilde{T}-\overline{T}\right)^{2}\nu^{a}\bm{\epsilon}_{abcd}.$ (100)
@@
$\delta_{2}\bm{J}^{\text{n}}|_{\mathcal{U}}$ $=-\delta_{2}\bm{J}^{\text{p}}|_{\mathcal{U}}=\varepsilon\kappa\left(\Omega_{\text{n}}-\Omega_{\text{p}}\right)\nu^{a}\bm{\epsilon}_{abcd},$
$\delta_{2}\bm{J}^{\text{n}}|_{\Sigma\backslash\mathcal{U}}$ $=\delta_{2}\bm{J}^{\text{p}}|_{\Sigma\backslash\mathcal{U}}=0,$ (101)
@@
$\delta_{2}N=\delta_{2}S=\delta_{2}J=0,$ (102)
@@
$\delta_{2}M$ $=\int_{\Sigma}\left(\sum_{\text{X}}\tilde{\mu}_{\text{X}}\delta_{2}\bm{N}^{\text{X}}+\tilde{T}\delta_{2}\bm{S}+\sum_{\text{X}}\Omega_{\text{X}}\delta_{2}\bm{J}^{\text{X}}\right)$
$=\varepsilon\int_{\mathcal{U}}\kappa\left(\Omega_{\text{n}}-\Omega_{\text{p}}\right)^{2}\nu^{a}\bm{\epsilon}_{abcd}.$ (103)
@@
And the equation of motion is given by [15]
@@
Now, in the transfusive case, without the conserved quantities given in Eq. (75), our definition of weak thermodynamic equilibrium obviously can not work, but the definition of strong thermodynamic equilibrium still works due to the conservation of the total particle number $N$, the total entropy $S$, and the total angular momentum $J$. As shown in [26], if there is a differential rotation, the different constituents will drag each other until they have same angular velocity. This statement actually coincides our condition that there will be no differential rotation as stated in Theorem 3. On the other hand, although the three fluids with differential rotation will finally be locked together, this progress likely takes many hundreds of dynamical timescales, so when the timescale under consideration is not too long, the transfusive effect is not important, and the non-transfusive model will be a good approximation to describe the neutron star.
@@
We want to construct the phase space of our Einstein-superconducting-superfluid system. To achieve this, we will give the pre-symplectic form $W$ by using Wald formalism introduced in Sec. III, and we will also determine the characteristic equation of sound velocity for our neutron star. With these preparations, we then seek the degeneracy of $W$ and construct the phase space by factoring these degeneracy. Moreover, since in the next section, it will be important to determine the symplectic complement corresponding to field variations of the trivial perturbation within the subspace consisting of weak solutions to the linearized constraints, so we will talk about the trivial perturbations generated by the trivial displacements in the last subsection.
@@ V.1 Lagrangian and symplectic form
And the constraint 3-form $\bm{C}_{\mathcal{X}}$ is still
$\left(\bm{C}_{\mathcal{X}}\right)_{abc}=-\mathcal{X}^{d}\bm{\epsilon}_{eabc}\left[2\left(E_{G}\right)_{\ d}^{e}+E_{F}^{e}A_{d}\right].$ (117)
@@
Since when the Lie derivative acting on the forms, one has $\left[\mathscr{L}_{\mathcal{X}},\mathscr{L}_{\mathcal{Y}}\right]=\mathscr{L}_{\left[\mathcal{X},\mathcal{Y}\right]}$, so that
@@
where we have used the well known expression [5] for the gravitational and electromagnetic parts of symplectic current, and
@@
V.2 Speed of sound and causal behavior
The purpose of this subsection is to determine the propagation speeds and polarization directions of sound wave fronts in our superconducting-superfluid star. A convenient method is the so-called Hadamard technique (introduced long ago by Hadamard [hadamard1903leçons] and then generalized to general-relativistic elasticity theory [Rayner:1963],[4]) of investigating the characteristic hypersurfaces of possible discontinuity in a partial differential system. By using such a method, Carter and Langlois succeeded in calculating the first and second sound in a relativistic two-constituent superfluid [13]. The method works by considering the first order case in which the algebraically related variables $n_{\text{X}}^{a}$, $s$, and $\mu_{a}^{\text{X}}$ are themselves continuous but have space-time derivatives that are weakly discontinuous across some characteristic hypersurface with tangent direction specified by some normal covector $\lambda_{a}$ say. For any component $\phi$, the discontinuities $\lceil\nabla_{a}\phi\rfloor$ in its gradient components will have to be proportional to the normal $\lambda_{a}$, i.e., we shall have $\lceil\nabla_{a}\phi\rfloor=\hat{\phi}\lambda_{a}$ for some scalar $\hat{\phi}$. Applying this to the relevant variables in the present case, we shall have
$\hat{n}_{\text{X}}^{a}\lambda_{a}$ $=0,\quad u_{\text{e}}^{a}\lambda_{a}\hat{s}=0,$
$\qquad\qquad+2\sum_{\text{Z}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial n_{\text{Y}}^{2}\partial x_{\text{XZ}}^{2}}n_{\text{Z}a}n_{\text{Y}b}+\sum_{\text{Z}\neq\text{Y}}\sum_{\text{W}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial x_{\text{YZ}}^{2}\partial x_{\text{XW}}^{2}}n_{\text{W}a}n_{\text{Z}b}\bigg)\delta n_{\text{Y}}^{b}$
$\qquad+\sum_{\text{Z}}\sum_{\text{Y}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial n_{\text{Z}}^{2}\partial x_{\text{XY}}^{2}}n_{\text{Y}a}n_{\text{Z}}^{b}n_{\text{Z}}^{c}+\sum_{\text{Z}\neq\text{W}}\sum_{\text{Y}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial x_{\text{WZ}}^{2}\partial x_{\text{XY}}^{2}}n_{\text{Y}a}n_{\text{Z}}^{b}n_{\text{W}}^{c}\bigg)\delta g_{bc}$
where the explicit expressions are
will be given for a suitably normalized $\lambda_{a}$ by
$\lambda_{a}=-vu_{\text{e}a}+\gamma_{a}.$ (130)
$vu_{\text{e}a}\hat{n}_{\text{X}}^{a}=\gamma_{a}\hat{n}_{\text{X}}^{a}.$ (133)
Consequently,
and
$2\gamma^{a}u_{\text{X}}^{b}\lambda_{[a}\hat{\mu}_{b]}^{\text{X}}$
The resulting eigenvalue equations for the propagation velocity $v$ is
V.3 Phase space
@@
$W\left(\delta_{1}\phi,\delta_{2}\phi\right)$
$=$ $W\left(\delta_{1}\phi,\delta_{2}^{\prime}\phi\right)+W\left(\delta_{1}\phi,\delta_{2}^{\text{M}}\phi\right)$
$+\int_{\Sigma}\left(\delta_{1}\bm{\Pi}^{i}\delta_{2}A_{i}-\delta_{2}\bm{\Pi}^{i}\delta_{1}A_{i}\right)-\int_{\Sigma}\sum_{\text{X}}\xi_{\text{X}}^{a}\delta_{2}^{\prime}\bm{P}_{a}^{\text{X}}$
@@
Note that the third and fourth lines only occur the quantities $\delta_{2}h_{ij}$, $\delta_{2}\bm{\pi}^{ij}$, $\delta_{2}\alpha$, $\delta_{2}\beta_{a}$, $\delta_{2}A_{a}$, $\delta_{2}\bm{\Pi}^{i}$, while the fifth line only occurs $\zeta_{\text{X}}^{a}$ and its derivative. According to
@@
where
@@
into Eq. (126) and $q_{\text{XY}}^{ab}=g^{ab}+\frac{1}{x_{\text{XY}}^{2}}n_{\text{X}}^{(a}n_{\text{Y}}^{b)}$. Also
@@
i.e., $\xi_{\text{X}}^{a}$ is the linear combination of $\nu^{a}$ and $\varphi^{a}$. Thus Eq. (163) amounts to say that the only choice is $\xi_{\text{X}}^{a}\propto u_{\text{X}}^{a}$. Substituting return to Eq. (155), one can easily see that $\xi_{\text{X}}^{a}\propto u_{\text{X}}^{a}$ is actually the solution. It immediately that Eqs. (148) and (149) are satisfied, in addition, Eqs. (150) and (153) yield $\delta_{1}\bm{\pi}^{ij}=0$ and $\delta_{1}\bm{\Pi}^{i}=0$ respectively.
@@
where we use $\cdot$” denotes the contraction of all tensor indices after using the background metric $h_{ab}$ on $\Sigma$ to raise and lower indices. Thus the elements of $\mathcal{H}$ are the square integrable tensor fields $\left(q^{\alpha},p_{\alpha}\right)$ on $\Sigma$. Note that the perturbations for which $\delta M\neq 0$ fall off too slowly to be square integrable, but $\mathcal{H}$ contains all perturbations of interest for which $\delta M=0$.
@@
Now let $\phi$ satisfy the equations of motion and let $\mathcal{X}^{a}$ be smooth and of compact support. By the fundamental identity Eq. (42), we have for all $\delta\phi\in\mathcal{H}$,
$\braket{\delta\phi,\hat{W}\mathscr{L}_{\mathcal{X}}\phi}=\int_{\Sigma}X^{a}\delta\bm{C}_{a}.$ (182)
By definition, the right side vanishes if and only if $\delta\phi$ is a weak solution of the constraint equations, $\delta\bm{C}_{a}=0$. Thus if we take $\mathcal{G}$ to be subspace of $\mathcal{H}$ spanned by perturbations of the form $\mathscr{L}_{\mathcal{X}}\phi$, we see that $\mathcal{G}^{\perp_{s}}$ is precisely the subspace, $\mathcal{C}$, of weak solutions to the linearized constraints. Furthermore, by the general argument of the previous paragraph, we have $\mathcal{C}^{\perp_{s}}=\overline{\mathcal{G}}$. Another way of saying this is that if we restrict the action of the original quadratic form to $\mathcal{C}\times\mathcal{C}$, it becomes degenerate precisely on the gauge transformations $\mathscr{L}_{\mathcal{X}}\phi$.
V.4 Trivial displacements